OpenStudy (anonymous):

limit as x approaches 2 from the left hand side for the following function [(x-27)(x-91)]/(x^2 + 4x + 4)

OpenStudy (anonymous):

\[\lim_{x \to 2^-} \frac{(x-27)(x-91)}{x^2+4x+4}\] then just plug in the value 2 and see where it gets you \[\lim_{x \to 2^-} \frac{(2-27)(2-91)}{2^2+4*2+4}\] \[\lim_{x \to 2^-} \frac{(-25)(-89)}{16}\] \[\lim_{x \to 2^-} \frac{2225}{16}\] \[\lim_{x \to 2^-} \frac{2225}{16} = 139.0625\]